OpenStudy (anonymous):

solve -81+4x^2=0 by factoring

OpenStudy (anonymous):

Remember the rule about factoring differences of squares: \(a^2-b^2 = (a+b)(a-b)\). Does that help you out?

OpenStudy (anonymous):

no can u explain it farther

OpenStudy (anonymous):

So are the two parts of the equation (\(81\) and \(4x^2\)) squares?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now set \(a=4x^2\) and \(b=81\), and plug it in to the equation I gave above.

OpenStudy (anonymous):

Uh... correction: \(a^2 = 4x^2\) and \(b^2 = 81\).

OpenStudy (anonymous):

Yep. (2x - 9) (2x + 9)

OpenStudy (anonymous):

\[-81+4x^2=0\]

OpenStudy (anonymous):

can u explain step by step because i am on these problems in my math class